OpenStudy (abstracthaven):

find slope intercept form and standard form throughh (-2,4) and (1,3). so far I have the slope (-1/3)using m=y2-y1/X2-X1 but Im not ure how to get the standard form

OpenStudy (anonymous):

If you take the slope and one of the points you can get the equation of a line in the form of y=mx+b. Set it up using y-y1=m(x-x1). I'll use (1,3) and m=-1/3. y-3=-1/3(x-1). Distribute the -1/3 to give you y-3=-x/3 +1/3 then transpose the -3 to the right side of the equation to end up with y=-x/3 + 8/3.

OpenStudy (abstracthaven):

so how do i get the standard form

OpenStudy (anonymous):

move the -x/3 back over to the left, so it looks like -x/3 + y = 8/3

OpenStudy (anonymous):

er, my bad, it is actually positive, sorry. x/3 +y = 8/3